OpenStudy (anonymous):

Find the limit lim (n>infinite) an where a1=4 and a(n+1)=5 - 1/an. i got L=(5+rt21)/2 or (5-rt21)/2 so which is the actual one??

OpenStudy (anonymous):

\[a_{n+1} =5-1/a_{n}\]

OpenStudy (anonymous):

a2=4.75 a3>a2 and a4>a3 so L>4 then L=(5+rt21)/2 is the correct answer

OpenStudy (anonymous):

aha, that's right!!1 i didnt recognize that ~~~ thx